# Open-Vocabulary Action Localization with Iterative Visual Prompting

### v4

@@ Abstract
Video action localization aims to find the timings of specific actions from a long video. Although existing learning-based approaches have been successful, they require annotating videos, which comes with a considerable labor cost. This paper proposes a learning-free, open-vocabulary approach based on emerging off-the-shelf vision-language models (VLMs). The challenge stems from the fact that VLMs are neither designed to process long videos nor tailored for finding actions. We overcome these problems by extending an iterative visual prompting technique. Specifically, we sample video frames and create a concatenated image with frame index labels, making a VLM guess a frame that is considered to be closest to the start and end of the action. Iterating this process by narrowing a sampling time window results in finding the specific frames corresponding to the start and end of an action. We demonstrate that this technique yields reasonable performance, achieving results comparable to state-of-the-art zero-shot action localization. These results illustrate a practical extension of VLMs for understanding videos. Sample code is available at https://microsoft.github.io/VLM-Video-Action-Localization/
@@ I Introduction
Video action localization aims to find the timings of specific actions from a long video. Extracting precise timings of an action is a fundamental technique in computer vision and related research domains, including video annotation, video editing, and automatic video collection from unlabeled datasets. For example, in the context of robotics, accurately extracting videos corresponding to specific actions will greatly benefit from analyzing human action in several robot-teaching frameworks where robots learn from human demonstrations, such as Learning from Demonstration (LfD)  [1] and Learning from Observation (LfO) [2, 3, 4].
@@
Previous mainstream research on video action localization has primarily focused on training models based on labeled datasets. The typical approach is a frame-by-frame method where a pipeline classifies frames or a series of consecutive frames with predefined action labels  [5, 6, 7]. While these methods have been successful on several benchmarks, they have limitations in label extensibility. Specifically, these methods require additional data collection and model retraining when new labels are added, which is a bottleneck for the quick and convenient application of these methods to data outside benchmark tasks.
In this paper, we introduce a pipeline for Open-vocabulary Video Action Localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called Prompting with Iterative Visual Optimization (PIVOT) [8]. The original PIVOT aims to utilize Vision-Language Models (VLMs) to reason an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find temporal segments of actions. Specifically, we sample video frames and create a concatenated image with frame index labels, making a VLM guess the frame closest to the start and end of an action. Iterating this process by narrowing the sampling time window results in finding specific frames when the action started and ended, thus achieving action localization. Although the recent advancements in VLM technology have supported the development of OVAL [9, 10, 11, 12, 13, 14], these methods still rely on domain-specific pre-training and to the best of our knowledge, there are no training-free methods that utilize off-the-shelf VLMs directly.
@@
Figure 1: Open-vocabulary temporal action localization aims to find the start and end timing of an action specified by open-vocabulary free-text queries. We propose a learning-free approach that leverages an off-the-shelf vision-language model (e.g., OpenAI’s GPT-4o).
@@
Figure 2: The proposed pipeline for open-vocabulary video action localization using a VLM consists of the following steps: (a) Frames are sampled at regular intervals from a time window, covering the entire video in the first iteration. (b) The sampled frames are then tiled in an image with annotations indicating the time order of the frames. (c) This image is then fed into a VLM to identify the frames closest to a specific timing of an action (e.g., the start timing of an action). (d) The sampling window is updated by centering on the selected frame with a narrower sampling interval. Bottom panel (1) For general action localization, the start time of the action in the video is determined by iterating steps (a) to (d). Bottom panel (2) By estimating the end time of the action in the same manner, the action is localized in the video.
@@
Using OpenAI’s GPT-4o [15] as an example of VLMs, we evaluated the performance of the proposed pipeline on an existing breakfast-cooking dataset and our own labeled cooking dataset. We checked how the pipeline works for a typical robot-teaching condition in which transitions between high-level tasks need to be grounded to a human demonstration video [16, 17, 18] (e.g., grasp, pick up, move, put, and release an object). As result, we show that the model achieved over 60% accuracy in terms of mean-over-frame (MoF), outperforming a baseline method. Although the performance appeared to be inferior to existing learning-based methods, the results demonstrated the feasibility of the proposed method. Furthermore, we tested the proposed pipeline for a general action-localization task using a sport video benchmark and showed that it performed comparably to state-of-the-art zero-shot methods. In addition, to offer further insights into its application, we explored the impact of several visual prompting parameters and present qualitative results for extracting specific actions from an unlabeled dataset. The contributions of this research are:
@@ II-A Model-based Action Localization
Model-based temporal action localization can be broadly categorized into anchor-based methods and boundary-based methods. Anchor-based methods, also known as top-down methods, mainly use multi-scale temporal sliding windows to detect pre-trained actions within untrimmed videos [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In contrast, boundary-based methods, also known as bottom-up methods, initially detect the temporal boundaries of actions, followed by further analysis of the proposal regions estimated from these boundaries [31, 6, 7, 27, 32, 33]. Boundary-based methods methods typically involve a two-stage pipeline consisting of temporal proposal generation and action classification [25, 27, 28, 26], while end-to-end approaches have also been proposed [34, 19, 35, 36]. The proposed method follows a boundary-based approach, as it aims to detect the start and end boundaries of actions. Additionally, unlike other methods, our VLM-based method accommodates open-vocabulary queries. This allows for the addition of new actions without the need for data collection or re-training.
II-B Open-vocabulary Action Localization
Open-vocabulary action localization can be broadly categorized into two approaches. The first approach outputs the closest matching action label from a pre-trained action set, while the latter is class-label-free, aiming to localize actions specified by a free-text query. In the former approach, LLMs or high-level scene understanding are used to bridge the gap between the query and pre-trained action labels [10, 9]. In a typical implementation of the latter approach, actions are first detected and then the corresponding action segment is determined based on the relationship between video features and the query [14]. Pre-trained multimodal models are often used to achieve these objectives [12, 13, 37, 38, 11]. The proposed method follows the class-label-free approach, leveraging off-the-shelf models without additional training. Its unique advantage lies in being training-free and using VLMs to directly match action detection with language in a single stage.
@@ III-A Core Pipeline
Figure 2 shows the core pipeline of the proposed method, aligning the original illustration of PIVOT [8]. The pipeline starts by sampling a number of frames at regular intervals from a time window of the given video (Fig. 2(a)). In the first iteration, the sampling window was set to be identical to the length of the video to capture the entire sequence of the video. The number of frames to sample is a hyperparameter and Fig. 2 illustrates the case of 16 frames. The sampled frames are then tiled in an image with the annotation of the time order of the frames (Fig. 2(b)). The image was then fed into a VLM to identify the frame closest to a specific action timing; Fig. 2(c) illustrates the start timing of an action. Finally, the sampling window is updated centering the selected frame with a narrower sampling interval (Fig. 2(d)). Note that due to the sparsity of sampling, the exact moment of an action may not necessarily be included in the query image. The pipeline can precisely identify the relevant moment by repeatedly narrowing the sampling window around the estimated closest frame until reaching a specific sampling interval. Hereafter, we will refer to this method as Temporal PIVOT (T-PIVOT).
For general action localization, we first estimate the start time of the action in the video using T-PIVOT (Fig. 2 bottom panel (1)). Then, by re-setting a sampling window for the period after the start time, we estimate the end time of the action in the same manner (Fig. 2 bottom panel (2)). By using the obtained start and end time, we localize the action of interest. Importantly, since this method involves querying the VLM, it allows us to specify actions in open-vocabulary free-text queries.
This pipeline can be extended to address several sub-scenarios, including but not limited to: finding the transitions between a series of actions within a video, detecting a specific event in a video, and collecting multiple instances of an action for a data-curating purpose.
@@ III-B Implementation of T-PIVOT
The Algorithm 1 shows the specific pipeline we implemented for the experiment in this paper. The computation yields the start and end times of a set of tasks in a video by applying T-PIVOT in parallel. We used OpenAI’s GPT-4o [15] for all the experiments. The sampling window is halved with each iteration, and this process is repeated for a specific number of iterations.
@@
  • $SampleFrames$ creates a grid of frames around $center\_time$ with an interval of $interval$ and returns the frames and their corresponding time.
  • $QueryVLM$ uses a VLM to analyze the image and determine a frame closest to the completion of the task.
@@ III-C Qualitative Results
We qualitatively checked our proposed pipeline (Fig. 2) using a cooking-preparation video that we recorded in-house. This 10-minute first-person video included actions such as taking out, washing, and cutting vegetables. Before applying Algorithm 1, the video was pre-divided into ten segments to maintain sufficient temporal resolution for recognition. We then extracted actions corresponding to open-vocabulary text descriptions. Figure 3 shows the examples of the identified video segments for actions of “cutting vegetables,” “washing vegetables,” and “turning on a faucet,” demonstrating that reasonable outputs were obtained.
Figure 3: Qualitative results of open-vocabulary action localization in a cooking-preparation video.
@@ IV Experiment
We evaluated the performance of the proposed pipeline in the light of practical use cases. Specifically, the pipeline was evaluated in the context of a typical robot-teaching condition on top of task and motion planning (TAMP)[39], where a system needs to ground high-level task plans (e.g., grasp, pick up, move, put, and release an object) to a human demonstration video [16, 17, 18]. To this end, we used an existing benchmark dataset and our own labeled cooking dataset. We focused on cooking actions because it reflects realistic scenarios as well as challenging appearance of cluttered objects in a scene.
@@ IV-A Benchmark
Breakfast Dataset [40] consists of third-person videos containing ten activities related to breakfast preparation, such as pouring a drink and making a sandwich. Each video is annotated without gaps along the timeline with more detailed action labels, such as taking a cup and stirring. The videos contain 52 different individuals across 18 different kitchens, totaling 1712 videos. As we are proposing a learning-free approach, we used only the testing dataset of 335 videos based on the split used in [41].
Fine-grained Breakfast dataset consists of the first-person videos of human cooking activities with fine-grained robotics tasks. Existing datasets of human actions including the Breakfast dataset [40] are labeled with coarse categories and simple motion labels often consist of multiple robotics manipulations. For example, a motion labeled as “pour milk” needs a sequential robotic manipulation of moving a robot hand towards a milk carton, grasp it, pick it up, move it near a cup, rotate and hold the milk carton, adjust its pose, move back to the table and release it [42, 43]—there is a huge gap between the granularity of action in the computer-vision community and the robotics community. To our knowledge, however, no existing datasets label tasks based on fine-grained robotic actions. To gain insights into the usefulness of the proposed pipeline for robot teaching, we prepared a fine-grained breakfast dataset by manually annotating an existing dataset of breakfast preparation [44] using a third-party annotation tool [45].
Based on the Kuhn-Tucker theory [46] explaining solution space of linear simultaneous inequalities in space, we defined tasks as transitions of constraints imposed on an object [47, 48, 49]. As such we obtain a task set that is sufficient and exhaustive to explain manipulation, such as picking up, moving, placing, rotating (refer to  [48] for the complete list of tasks). 113 videos were used in this study.
@@ IV-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Section III, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), is estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), is similarly estimated. This approach yields an estimated series of start and end times for each task. In this experiment, we assume that tasks are continuously present in the video. Therefore, the end time of one task is identical to the start time of the subsequent task. To ensure this, we calculate this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@
Specifically, the transition time $T_{i\rightarrow i+1}$ ($i=1\ldots N-1$) is calculated using Algorithm 2.
@@
$EstimateStartTimes$ corresponds to Algorithm 1

### v5

@@ Abstract
Video action localization aims to find the timings of specific actions from a long video. Although existing learning-based approaches have been successful, they require annotating videos, which comes with a considerable labor cost. This paper proposes a training-free, open-vocabulary approach based on emerging off-the-shelf vision-language models (VLMs). The challenge stems from the fact that VLMs are neither designed to process long videos nor tailored for finding actions. We overcome these problems by extending an iterative visual prompting technique. Specifically, we sample video frames and create a concatenated image with frame index labels, allowing a VLM to identify the frames that most likely correspond to the start and end of the action. By iteratively narrowing the sampling window around the selected frames, the estimation gradually converges to more precise temporal boundaries. We demonstrate that this technique yields reasonable performance, achieving results comparable to state-of-the-art zero-shot action localization. These results support the use of VLMs as a practical tool for understanding videos. Sample code is available at https://microsoft.github.io/VLM-Video-Action-Localization/
@@ I Introduction
Video action localization aims to find the timings of specific actions from a long video. Extracting precise timing of an action is a fundamental technique in computer vision and related research domains, including video annotation, video editing, and automatic video collection from unlabeled datasets. For example, in the context of robotics, accurately extracting videos corresponding to specific actions will greatly benefit from analyzing human action in several robot-teaching frameworks where robots learn from human demonstrations, such as Learning from Demonstration (LfD) [1] and Learning from Observation (LfO) [2, 3, 4].
@@
Previous mainstream research on video action localization has primarily focused on training models based on labeled datasets. The typical approach is a frame-by-frame method where a pipeline classifies frames or a series of consecutive frames with predefined action labels  [5, 6, 7]. While these methods have been successful on several benchmarks, they have limitations in label extensibility. Specifically, these methods require additional data collection and model re-training when new labels are added, which is a bottleneck for the quick and convenient application of these methods to data outside benchmark tasks.
@@
Figure 1: Open-vocabulary video action localization aims to find the start and end times of an action specified by open-vocabulary free-text queries. We propose a training-free approach that leverages an off-the-shelf vision-language model (e.g., OpenAI’s GPT-4o).
@@
Figure 2: The proposed pipeline for open-vocabulary video action localization using a VLM consists of the following steps: (a) Frames are sampled at regular intervals from a time window, covering the entire video during the first iteration. (b) The sampled frames are then tiled in an image with annotations indicating the time order of the frames. (c) This image is then fed into a VLM to identify the frames closest to a specific timing of an action (e.g., the start time of an action). (d) The sampling window is updated by centering on the selected frame with a narrower sampling interval. Bottom panel (1) For general action localization, the start time of the action in the video is determined by iterating steps (a) to (d). Bottom panel (2) By estimating the end time of the action in the same manner, the action is localized in the video.
In this paper, we introduce a pipeline for Open-vocabulary Video Action Localization (OVAL), which does not require pre-training and can address action labels in an open-vocabulary form (Fig. 1). Our approach is inspired by a recently proposed method called Prompting with Iterative Visual Optimization (PIVOT) [8]. The original PIVOT aims to utilize Vision-Language Models (VLMs) to infer an embodied action by framing the problem as iterative visual question answering. In this paper, we propose a simple yet effective approach to apply PIVOT to find temporal segments of actions. Specifically, we sample video frames and create a concatenated image with frame index labels, allowing a VLM to identify the frames that most likely correspond to the start and end of the action. By iteratively narrowing the sampling window around the selected frames, the estimation gradually converges to more precise temporal boundaries, thereby achieving action localization. Although the recent advancements in VLM technology have supported the development of OVAL [9, 10, 11, 12, 13, 14], these methods still rely on domain-specific pre-training, and to the best of our knowledge, there are no training-free methods that utilize off-the-shelf VLMs directly.
@@
Using OpenAI’s GPT-4o [15] as an example of VLMs, we evaluated the performance of the proposed pipeline on an existing breakfast cooking dataset and our own labeled cooking dataset. We checked how the pipeline works for a typical robot-teaching condition in which transitions between high-level tasks need to be grounded to a human demonstration video [16, 17, 18] (e.g., grasp, pick up, move, put, and release an object). As a result, we show that the model achieved over 60% accuracy in terms of mean-over-frame (MoF), outperforming a baseline method. Although the performance appeared to be inferior to existing learning-based methods, the results demonstrated the feasibility of the proposed method. Furthermore, we tested the proposed pipeline for a general action-localization task using a sport video benchmark and showed that it performed comparably to the state-of-the-art zero-shot methods. In addition, to offer further insights into its application, we explored the impact of several visual prompting parameters and present qualitative results for extracting specific actions from an unlabeled dataset. The contributions of this research are:
@@ II-A Model-based Action Localization
Model-based temporal action localization can be broadly categorized into anchor-based methods and boundary-based methods. Anchor-based methods, also known as top-down methods, mainly use multi-scale temporal sliding windows to detect pre-trained actions within untrimmed videos [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. In contrast, boundary-based methods, also known as bottom-up methods, initially detect the temporal boundaries of actions, followed by further analysis of the proposal regions estimated from these boundaries [31, 6, 7, 27, 32, 33]. Boundary-based methods typically involve a two-stage pipeline consisting of temporal proposal generation and action classification [25, 27, 28, 26], while end-to-end approaches have also been proposed [34, 19, 35, 36]. The proposed method follows a boundary-based approach, as it aims to detect the start and end boundaries of actions. Additionally, unlike other methods, our VLM-based method accommodates open-vocabulary queries. This allows for the addition of new actions without the need for data collection or re-training.
II-B Open-vocabulary Video Action Localization
OVAL can be broadly categorized into two approaches. The first approach outputs the closest matching action label from a pre-trained action set, while the latter is class-label-free, aiming to localize actions specified by a free-text query. In the former approach, LLMs or high-level scene understanding are used to bridge the gap between the query and pre-trained action labels [10, 9]. In a typical implementation of the latter approach, actions are first detected and then the corresponding action segment is determined based on the relationship between video features and the query [14]. Pre-trained multimodal models are often used to achieve these objectives [12, 13, 37, 38, 11]. The proposed method follows the class-label-free approach, leveraging off-the-shelf models without additional training. Its unique advantage lies in being training-free and using VLMs to directly match action detection with language in a single stage.
Recent studies have also highlighted the significance of open-vocabulary approaches in various other tasks beyond action localization. For instance, open-vocabulary learning has been extensively surveyed in [39], and open-vocabulary methods have been proposed for multiple object tracking [40], segmentation [41], and video instance segmentation [42], demonstrating a growing interest in more generalized and flexible VLMs. These approaches share a common goal of avoiding the close-set assumption, thereby offering broader applicability and robustness for real-world tasks.
@@ III-A Core Pipeline
Figure 2 shows the core pipeline of the proposed method, aligning with the original illustration of PIVOT [8]. The pipeline starts by sampling a number of frames at regular intervals from a time window of the given video (Fig. 2(a)). In the first iteration, the sampling window is set to be identical to the length of the video to capture the entire sequence of the video. The number of frames to sample is a hyperparameter and Fig. 2 illustrates the case of 16 frames. The sampled frames are then tiled in an image with the annotation of the time order of the frames (Fig. 2(b)). The image is then fed into a VLM to identify the frame closest to a specific action timing; Fig. 2(c) illustrates the start time of an action. Finally, the sampling window is updated by centering on the selected frame with a narrower sampling interval (Fig. 2(d)). Note that due to the sparsity of sampling, the exact moment of an action may not be present in the query image during the first iteration. However, the method can still identify the relevant moment precisely by iteratively narrowing the sampling window around the most likely frame until a specific temporal resolution is reached. Hereinafter, we will refer to this method as Temporal PIVOT (T-PIVOT).
For general action localization, we first estimate the start time of the action in the video using T-PIVOT (Fig. 2 bottom panel(1)). Then, by resetting a sampling window for the period after the start time, we estimate the end time of the action in the same manner (Fig. 2 bottom panel(2)). By using the obtained start and end times, we localize the action of interest. Importantly, since this method involves querying the VLM, it allows us to specify actions in open-vocabulary free-text queries.
This pipeline can be extended to address several sub-scenarios, including but not limited to: finding the transitions between a series of actions within a video, detecting a specific event in a video, and collecting multiple instances of an action for the purpose of data curation.
@@ III-B Implementation of T-PIVOT
Algorithm 1 shows the specific pipeline we implemented for the experiment in this paper. The computation yields the start and end times of a set of tasks in a video by applying T-PIVOT in parallel. We used GPT-4o for all the experiments. The sampling window was halved with each iteration, and this process was repeated for a specific number of iterations.
@@
  • $SampleFrames$ samples a set of frames centered at $center\_time$ within the given $window$, arranges them into a grid image, and returns the resulting image.
  • $QueryVLM$ uses a VLM to analyze the image and returns the $selected\_time$ corresponding to the frame that most likely represents the target moment of a given $task$.
@@ III-C Qualitative Results
We qualitatively checked our proposed pipeline using a cooking video that we recorded in-house (Fig. 2). This 10-minute first-person video included actions such as taking out, washing, and cutting vegetables. To maintain sufficient temporal resolution, we divided the video into five-second intervals and applied Algorithm 1 to each interval. If an action was detected, the subsequent window was adjusted to begin at the estimated end time of the localized action; otherwise, the window advanced by another five seconds. Figure 3 shows the examples of video segments identified for specified actions, demonstrating that the method produced reasonable results.
Notably, T-PIVOT worked not only for static actions that are closely tied to stationary scene information but also for dynamic hand movements. For instance, actions such as “cutting vegetables,” “washing vegetables,” and “turning on a faucet” could be recognized from a single frame due to the presence of static objects like a cutting board or faucet. In contrast, actions such as “picking something up” or “passing something from one hand to the other” would not rely on such contextual cues. Instead, they require attention to dynamic hand motion for accurate localization, which cannot be achieved solely by single-frame analysis. Nevertheless, our method successfully localized these actions based on textual descriptions. This success is likely because the input image, formatted by arranging time-series frames, conveyed temporal information such as hand motions to the VLM.
Another notable implication is that T-PIVOT localized second-scale actions within a long, 10-minute video. In cases where the target action occupies only a small portion of the video, the initial iteration of T-PIVOT is unlikely to extract the relevant frames, thus VLMs will not select the frames that include the action. However, by defining a reasonable temporal window (i.e., five seconds) and applying T-PIVOT sequentially, T-PIVOT effectively extracts actions regardless of the original video length. This temporal-window strategy was also employed in the experiment in zero-shot action localization in Sec. IV-F.
Figure 3: Qualitative results of open-vocabulary video action localization on a cooking video.
@@ IV Experiment
We evaluated the performance of the proposed pipeline in light of practical use cases. Specifically, the pipeline was evaluated in the context of a typical robot-teaching condition on top of task and motion planning (TAMP) [43], where a system needs to ground high-level task plans (e.g., grasp, pick up, move, put, and release an object) to a human demonstration video [16, 17, 18]. To this end, we used an existing benchmark dataset and our own labeled cooking dataset. We focused on cooking actions because they reflect realistic scenarios and pose challenges due to the presence of cluttered objects in the scene.
@@ IV-A Benchmark
Breakfast Dataset [44] consists of third-person videos containing ten activities related to breakfast preparation, such as pouring a drink and making a sandwich. Each video is annotated without gaps along the timeline with more detailed action labels, such as taking a cup and stirring. The videos contain 52 different individuals across 18 different kitchens, totaling 1712 videos. Since we are proposing a training-free approach, we used only the testing dataset of 335 videos based on the split used in [45].
Fine-grained Breakfast Dataset consists of first-person videos of human cooking activities with fine-grained robotics tasks. Existing datasets of human actions, including the Breakfast Dataset [44], are labeled with coarse categories and simple motion labels often consist of multiple robotic manipulations. For example, a motion labeled as “pour milk” needs a sequential robotic manipulation of moving a robot hand towards a milk carton, grasping it, picking it up, moving it near a cup, rotating and holding the milk carton, adjusting its pose, moving back to the table, and releasing it [46, 47]—there is a huge gap between the granularity of action in the computer vision community and the robotics community. To our knowledge, however, no existing datasets label tasks based on fine-grained robotic actions. To gain insights into the usefulness of the proposed pipeline for robot teaching, we prepared a Fine-grained Breakfast Dataset by manually annotating an existing dataset of breakfast preparation [48] using a third-party annotation tool [49].
Based on the Kuhn-Tucker theory [50] explaining solution space of linear simultaneous inequalities, we defined tasks as transitions of constraints imposed on an object [51, 52]. As a result, we obtained a task set that is necessary and sufficient to explain manipulation, such as picking up, moving, placing, and rotating (refer to [51] for the complete list of tasks). 113 videos were used in this study.
@@ IV-C T-PIVOT for Recognizing Seamless Action Transitions
By applying the T-PIVOT method described in Sec. III, we estimated the timing of task transitions within the video as follows. First, the start time of each task, $T_{i\_start}$ ($i=1\ldots N$), was estimated using the T-PIVOT in parallel processing. The end time of each task, $T_{i\_end}$ ($i=1\ldots N$), was similarly estimated. This approach yielded an estimated series of start and end times for each task. In this experiment, we assumed that tasks were seamlessly connected throughout the video. Therefore, the end time of one task was identical to the start time of the subsequent task. To ensure this, we calculated this time as the midpoint between the estimated end time of the current task and the estimated start time of the next task:
@@
Specifically, the transition time $T_{i\rightarrow i+1}$ ($i=1\ldots N-1$) was calculated using Algorithm 2.
@@
$EstimateStartTimes$ and $EstimateEndTimes$ correspond to Algorithm 1, which aims to find the start and end times, respectively.